OpenStudy (anonymous):

Which of the following can not be used on a matrix when solving a system of equations? interchange rows multiply a row by a constant add a constant to a row

OpenStudy (anonymous):

The third one

OpenStudy (anonymous):

Example: \[(1)\begin{cases}x+y=1\\ x-y=2\end{cases}\] Interchanging the rows gives an identical system: \[\begin{cases}x-y=2\\ x+y=1\end{cases}\] Multiplying a row by a constant doesn't fundamentally change the system; you can always undo that operation: \[\begin{cases}2x+2y=2\\ x-y=2\end{cases}\] Adding a constant to a row makes a completely new system: \[(2)\begin{cases}x+y=2\\ x-y=2\end{cases}\] System (1) has solution \(x=\dfrac{3}{2},y=-\dfrac{1}{2}\) whereas system (2) has solution \(x=2,y=0\).